A stamp collection consists of 5, 13, and 18 cent stamps. The number of 13 cent stamps is one less than twice the number of 5 cent stamps. The number of 18 cent stamps is five more than the number of 13 cent stamps. The total value of all the stamps is $2.60. Find the number of 13 cent stamps in the collection.
This is what I had tried to figure out
5 cent stamp = X
13 cent stamp = 2x-1
18 cent stamp = 2x-1+5
x=2x-1=2x-1+5 = 2.60

Answer by stanbon(75887)

A stamp collection consists of 5, 13, and 18 cent stamps.
The number of 13 cent stamps is one less than twice the number of 5 cent stamps.
Let # of 5 cent stamps be "x: Value is 5x cents
# of 13 cent stamps is 2x-1; value is 13(2x-1)=26x-13 cents
-----------------------
The number of 18 cent stamps is five more than the number of 13 cent stamps.
# of 18 cent stams is 2x-1+5 = 2x+4; value is 18(2x+4)= 36x+72 cents
------------------------
The total value of all the stamps is $2.60.
EQUATION:
value + value + value = 260 cents
5x+26x-13+36x+72 = 260
67x +59 = 260
67x = 201
x = 3
------------------------------------
Find the number of 13 cent stamps in the collection.
26x-13 = 26*3-13 = 65 (# of 13-cent stamps)
====================
Cheers,
Stan H.